OpenStudy (anonymous):

Express the integral as a limit of sums. int_{0}^{π}sin5xdx

OpenStudy (anonymous):

It is just 2/5

OpenStudy (amistre64):

int = sum; dx = \(\Delta x\)

OpenStudy (amistre64):

maybe?

OpenStudy (anonymous):

oh, you mean like approximation; midpoint, simpson..

OpenStudy (amistre64):

limit of sums reminds me of reimann stuff

OpenStudy (jamesj):

It comes back to the partition question you asked earlier. The Riemann integral is defined to be the limit of a Riemann sum. Here's such a sum where the interval [0,pi] has just two partitions, [0,pi/2], [pi/2,pi]. I.e., we've broken up the interval of integration into two intervals, each of length pi/2: \[ S_n = \frac{\pi}{2} \sin(0) + \frac{\pi}{2} \sin(\pi/2) \]

OpenStudy (jamesj):

That is in fact such a sum \( S_2 \), 2 for two partitions of the interval [0,pi].

OpenStudy (amistre64):

\[\lim_{\Delta x ->0}\ \sum_{i=0}^{\pi} f(x_i)\Delta x_i\]

OpenStudy (jamesj):

You can see that it's an approximation for the area, or the integral. It's a bad approximation, but that's to be expected for a partition with such a small number.

OpenStudy (jamesj):

If we broke [0,pi] up into 4 partitions of equal length pi/4, the Riemann sum would be \[ S_4 = \frac{\pi}{4} \sin(0) + \frac{\pi}{4} \sin(\pi/4) + \frac{\pi}{4} \sin(2\pi/4) + \frac{\pi}{4} \sin(3\pi/4) \]

OpenStudy (jamesj):

This is a better approximation of the integral. Following?

OpenStudy (anonymous):

yes, sir!

OpenStudy (jamesj):

So now we can generalize to n partitions of [0,pi], in which case the Riemann sum is \[ S_n = \sum_{i=0}^{n-1} \sin(0 + i \frac{\pi}{n}) .\frac{\pi}{n} \] I wrote the zero in the sin just to emphasize that we're beginning at zero.

OpenStudy (jamesj):

If now the limit of the Riemann sum S_n exists as n --> infinity, then the Riemann integral over that interval exists and we write \[ \lim_{n \rightarrow \infty} S_n = \int_0^{\pi} \sin x \ dx \] The definition of the symbols on the right hand side of the equation is the limit.

OpenStudy (jamesj):

(There are some minor technical wrinkles here, but this is substantially correct.)

OpenStudy (jamesj):

And for your problem you need to use sin(5x), not sin(x) as I have.

OpenStudy (anonymous):

yeah, I got that... Thanks :)